OpenStudy (anonymous):

Problem Set 1 Question 1A-1 b) By completing the square, use translation and change of scale to sketch: y = 3x^2 + 6x +2 Here is the method I attempted, but with a different result than shown in the Solutions. Where am I going wrong? y = 3x^2 + 6x +2 y - 2 = 3x^2 + 6x [move 2 to the left side] y - 2 = 3(x^2 +2x) [factor out 3] y - 2 +1 = 3(x^2 +2x +1) [complete the square] y - 1 = 3(x +1)^2 [rearrange toward slope - intercept form] y = 3(x+1)^2 +1 [add one to both sides] Now, I have it in slope-intercept form but my answer is different from the Solution. Help?

OpenStudy (anonymous):

In your fourth step, you've added 1 to the left side, but 3 to the right (3(+1), so the quantities are no longer equal. Try adding 3 to both sides, and you should get what you're looking for :)

OpenStudy (anonymous):

\[y-2=3(x ^{2}+2x)\] Next you'll add 3 to both sides: \[y-2+3=3(x ^{2}+2x)+3\] Factor out the 3, \[y+1=3(x ^{2}+2x+1)\] and take it from there.

OpenStudy (anonymous):

Thank you, Throol. I see that your solution is equivalent to the original equation.